# arXiv:2603.00175v1[cs.CV] 26 Feb 2026

### v1

@@ 4.2 Attention Quality: Degradation and Localization
Figure 11: MoRF and LeRF curves comparison. Combined MoRF degradation and LeRF retention curves for all attention mechanisms. InfSA variants exhibit the sharpest MoRF degradation and the highest LeRF retention, confirming that they assign both high attention to salient regions and low attention to irrelevant background.
Figure 12: Combined MoRF–LeRF summary. Joint view of MoRF and LeRF metrics across attention mechanisms. InfSA variants achieve the best trade-off between foreground sensitivity (MoRF) and background rejection (LeRF).
Fig. 11 overlays the MoRF degradation and LeRF retention curves for all three attention mechanisms on a single plot. The ideal attention map would produce a steep MoRF curve (removing high-attention patches quickly destroys classification confidence) and a high LeRF curve (removing low-attention patches preserves confidence, because they correspond to background). InfSA variants excel on both axes simultaneously: Pure InfSA achieves the highest LeRF retention (65.3%) while also delivering strong MoRF degradation (71.7%), and Linear InfSA leads MoRF (76.0%) with competitive LeRF (64.0%). Standard ViT’s curves are nearly flat for MoRF (42.6% AOC) and lower for LeRF (60.1%), indicating that its attention spreads indiscriminately across foreground and background patches.
Fig. 12 presents the same trade-off as a joint summary plot, making it easy to compare mechanisms at a glance. The vertical gap between InfSA and Standard ViT is 28–33 pp on MoRF-AOC and 4–5 pp on LeRF-AUC—substantial, consistent improvements that hold across both the patch-masking (MoRF/LeRF) and the bounding-box (ROC/PR) evaluation families.
4.2.2 Bounding-Box Attention Localization.
Figure 13: Attention bounding-box localization. Overlay of attention maps with ImageNet ground-truth bounding boxes. InfSA variants concentrate attention mass inside object regions, while Standard ViT distributes attention diffusely across the image.
Figure 14: Bounding-box localization comparison. Per-method localization accuracy using attention as a detector of ImageNet bounding-box regions. InfSA variants achieve substantially higher localization precision than Standard ViT and sub-quadratic baselines.
Figs. 13 and 14 visualize and quantify attention localization against ImageNet ground-truth bounding boxes (2 088 images that have bounding-box annotations in the validation set).
Fig. 13 shows representative overlays of attention heatmaps on input images with the ground-truth bounding box drawn in white. For Standard ViT, attention mass is spread across background regions and often peaks outside the bounding box. Both InfSA variants concentrate the heat inside the object boundary, producing sharper and more object-aligned maps. This visual difference is consistent across diverse object categories and scales.
Fig. 14 translates the visual impression into a per-method quantitative comparison, plotting localization accuracy (the fraction of total attention mass falling inside the bounding box) for each mechanism. Pure InfSA leads in ROC-AUC (77.3%) and Linear InfSA in PR-AUC (76.1%), versus 53.8% and 56.2% for Standard ViT—improvements of ++23 pp and ++20 pp, respectively. These results are important because they demonstrate that InfSA’s attention is not merely “different” from softmax attention but more faithful to the semantic content of the image, as independently verified by ground-truth object annotations.
4.3 ImageNet Classification Results
All models are trained on ImageNet-1K (1.28M images, 300 epochs, batch 64) with a DeiT-style recipe [58], without external data, distillation, or self-supervised pretraining. Our ViT baseline reaches 81.5%; the +3.2+3.2 pp gain of Linear InfViT 4L (84.7%, 53.5M params, 59 GFLOPs) is purely architectural. Note that InfViT-4L operates at 59 GFLOPs ($224^{2}$) vs. 17.5 GFLOPs for ViT-B/16, owing to its 64-head design. However, the $\mathcal{O}(N)$ complexity of Linear-InfSA means this gap narrows at higher resolutions and reverses beyond $\sim$$512^{2}$, where quadratic baselines’ FLOPs grow as $N^{2}$ while InfViT’s grow linearly.
Fig. 15 plots ImageNet-1K and ImageNet-V2 [44] top-1 accuracy against parameter count. On IN-1K, InfViT Pure 24L reaches 85.4%, within 0.4 pp of RAVLT-L [15] (85.8%). The 4L variants (85.1% Pure, 84.7% Linear) surpass Agent Attn-B (84.1%) and FLatten ($\leq$}84.5%) at roughly half the parameters. On IN-V2, all four InfViT models exceed every baseline (up to 79.8% vs. 76.8% for RAVLT-L), indicating robust generalization. The 24L models improve by only 0.3–0.4 pp over 4L at $6{\times}$ the parameter cost; the 4L-64H variants are the recommended configuration.
4.3.1 Full Classification Table.
Table 3 reports per-method Top-1 accuracy on ImageNet-1K and ImageNet-V2, grouped by attention family.
@@ 4.3 ImageNet Classification Results
ViT-B/16 [13] ’21 86 17.6 81.5 74.1
DeiT-B [58] ’21 86 17.5 81.8 71.5
HorNet-B [43] ’22 88 15.5 84.3
SG-Former-B [45] ’23 78 15.6 84.7
SMT-L [36] ’23 81 17.7 84.6
InternImage-B [63] ’23 97 16.0 84.9
GC-ViT-B [25] ’23 90 14.8 85.0
FAT-B5 [17] ’23 88 15.1 85.2
STViT-L [28] ’23 95 15.6 85.3
RMT-L [16] ’24 95 18.2 85.5 76.3
@@
SOFT-H [38] ’21 87 16.3 83.3
FLatten-S-B [21] ’23 89 15.4 83.8
FLatten-C-B [21] ’23 75 15.0 84.5
ViG-S [35] ’24 23 3.5 83.4
Agent Attn-B [23] ’24 88 15.4 84.1
MLLA-B [22] ’24 96 16.2 85.3 76.7
RAVLT-L [15] ’24 95 16.0 85.8 76.8
@@
HyenaPixel [54] ’24 111 25.3 84.9
MambaVision-L [24] ’25 228 34.9 85.0
MambaVision-L2 [24] ’25 242 37.5 85.3
@@
InfViT-Linear (4L) Ours 53.5 59.0 84.7 77.4
InfViT-Pure (4L) Ours 57.7 59.0 85.1 79.2
InfViT-Linear (24L) Ours 305 85.1 77.7
InfViT-Pure (24L) Ours 331 85.4 79.8
@@ 4.4 Ablation Study
Key observations.
(i) Path-decay $\gamma$. Top-1 accuracy and MoRF-AOC both peak at $\gamma{=}0.7$. Smaller values ($\gamma{=}0.3,0.5$) under-weight multi-hop paths, limiting the effective receptive field of the Neumann series; larger values ($\gamma{=}0.9$) over-weight distant interactions and introduce noise from long, weakly relevant paths. The optimum at $\gamma{=}0.7$ balances local precision and global context.
(ii) LeRF-AUC increases monotonically with $\gamma$ (62.1% $\to$ 65.3%), indicating that deeper path propagation helps the model assign low attention to irrelevant background regions. This is consistent with the graph-diffusion interpretation: a larger $\gamma$ extends the effective random-walk horizon, allowing the model to “see through” intermediate patches and concentrate on structurally central tokens.
(iii) Latency and convergence are stable across all settings (35–37 ms train, 11–12 ms infer, 40–45 epochs to convergence), confirming that changing $\gamma$ or the activation function does not alter the computational graph’s structure, only its numerical values. This is expected because $\gamma$ only scales the attention weights—it does not add operations or change tensor shapes.
(iv) Activation function. Among activations, absolute value $|x|$ yields the highest LeRF-AUC (65.5%) and ROC-AUC (77.2%), suggesting that preserving the sign of negative attention scores (via $|x|$ instead of clamping to zero with ReLU) retains additional background-awareness. However, $|x|$ reduces Top-1 accuracy by 1.9 pp relative to ReLU, likely because the non-selective activation does not suppress noisy token interactions during training. GELU, which is the standard activation in modern Transformers, performs worst on all attention-quality metrics, probably because its smooth saturation makes the power-iteration-like eigenvector approximation less sharp. ReLU offers the best overall trade-off between classification accuracy (84.7%) and attention quality (MoRF 76.0%, PR-AUC 76.6%) and is adopted as default ($\gamma{=}0.7$, ReLU).
4.5 Eigenvector Alignment Validation
Experimental setup.
Ground-truth Perron eigenvector.
Reconstruction of Linear-InfSA weights.
From the same tensors, we reconstruct the weight vector $a$ using the closed-form equations from Sec. 3.4:
$a_{t}$ $=\frac{s_{t}}{\textstyle\sum_{s}s_{s}+\varepsilon}.$ (27)
Metrics.
Metric Mean Std. dev.
Cosine similarity $0.985$ $0.036$
Spearman correlation $0.937$ $0.132$
Discussion.
Table 5 shows that the Linear-InfSA weights are almost perfectly aligned with the Perron eigenvector: mean cosine 0.985 ($\pm$0.036) and mean Spearman 0.937 ($\pm$0.132). The high cosine similarity means that the closed-form weights $a$ and the power-iteration eigenvector $v$ point in nearly the same direction in $\mathbb{R}^{T}$, and the high Spearman correlation means they induce nearly the same ranking of tokens—which is what matters for attention weighting. The slightly larger variance of the Spearman metric is expected because rank correlation is more sensitive to perturbations among tokens with similar centrality scores.
This result validates the theoretical claim that Linear-InfSA faithfully recovers the dominant spectral structure of the full $N{\times}N$ operator $\hat{A}=\phi(QQ^{\top})$ without ever forming it. The practical implication is that Linear-InfSA’s $\mathcal{O}(N)$ cost comes with negligible approximation error: the 1.5% gap in cosine similarity corresponds to roughly 0.4 pp in classification accuracy (84.7% Linear vs. 85.1% Pure at 4L), a small price for a $13{\times}$ speed-up and the ability to scale to $9216^{2}$ resolution.
4.6 Qualitative Analysis of Attention Maps
@@
[…]
Semantic localization. InfSA variants (both Pure and Linear) exhibit sharper focus on semantically meaningful object regions (faces, limbs, object centres), whereas softmax attention often highlights diffuse or peripheral areas. For example, on the “goldfish” image (top row), softmax attention distributes weight across the water background, while both InfSA variants precisely outline the fish body. On the “dog” image, InfSA concentrates on the head and torso, whereas softmax attention leaks into the lawn. These patterns are consistent across diverse object categories and scales.
Consistency across categories. Both InfSA variants produce attention distributions that are tighter and more consistent across samples and categories, suggesting higher robustness to irrelevant context. We attribute this to the spectral nature of InfSA: by modeling attention as a power series over token interactions (Pure InfSA) or approximating the dominant eigenvector of the token graph (Linear-InfSA), the mechanism recursively amplifies the most structurally central tokens and suppresses peripheral ones. This “rich get richer” dynamic—well-studied in graph-centrality theory [5, 6]—explains why InfSA attention maps are sharper than softmax even without any explicit sparsification or top-$k$ masking. The Markov-chain interpretation further explains the consistency: the fundamental-matrix entry $N_{ij}$ is a global property of the graph, less sensitive to local noise than the pointwise softmax score $\exp(q_{i}^{\top}k_{j}/\sqrt{d})$.
4.7 Reproducibility and Implementation Details
4.7.1 Latency vs. Resolution.
@@ 4.7 Reproducibility and Implementation Details
Fig. 17 shows latency as a function of input resolution from $224^{2}$ to $4096^{2}$ (where all models still fit in memory on the RTX 5090 32 GB used for this particular measurement). Dotted lines indicate training latency; solid lines indicate inference. Linear InfViT’s curves are nearly straight on the log-log axes—the hallmark of linear scaling—while the softmax ViT curves bend upward quadratically. The slope separation between the two becomes clearly visible above $512^{2}$, confirming the theoretical crossover point discussed in Sec. 4.3. Note: this figure uses an RTX 5090 (32 GB); all other benchmarks in the paper use an A100 40 GB. Relative scaling trends are consistent across GPUs.
Figure 18: Training vs. inference latency across resolutions (A100 40 GB). Full comparison of training (filled markers) and inference (open markers) latency for all nine attention mechanisms from $224^{2}$ to each method’s maximum resolution before OOM. Linear InfViT maintains the lowest latency at every resolution tested and is the only method that completes $4096^{2}$ training and $9216^{2}$ inference.
Fig. 18 provides the full training-vs-inference latency comparison on the A100 benchmark GPU. At $224^{2}$, the latency differences are modest (all methods complete in 4–60 ms), but they diverge rapidly: at $1024^{2}$, Standard ViT’s inference latency reaches 113 ms while InfViT Linear remains at 4.3 ms. Beyond $1024^{2}$, all non-InfViT methods trigger OOM errors (their data points stop), while InfViT Linear continues to $9216^{2}$ at $\sim$320 ms—demonstrating both the asymptotic advantage and the practical resilience of the $\mathcal{O}(N)$ formulation.
4.7.2 Hardware and Software Environment.
@@
4.7.5 Model Architecture and FLOPs.
@@
4.7.6 Extreme-Resolution Inference.
@@ 5 Discussion and Future Directions
The spectral and diffusion viewpoints rely on standard assumptions (homogeneous operators in the Neumann-series discussion, nonnegativity in Perron–Frobenius analysis). These serve as modelling tools for design motivation; our eigenvector-alignment study (Sect. 4.5) confirms they hold in practice for trained models.
@@
6 Conclusion
Empirically, InfViT models show strong performance across classification, localization, and scalability tasks. Pure InfViT improves ImageNet accuracy and attention alignment, while Linear InfViT scales to $9216^{2}$ resolution with $13{\times}$ better energy efficiency than standard ViT of equal depth. Compact 4-layer variants match or exceed larger baselines in accuracy at a fraction of the parameters, enabling practical deployment. The graph-theoretic principles underlying InfSA are modality-agnostic, suggesting natural extensions to NLP, multi-modal models, video understanding, and dense prediction tasks. Code and pretrained weights are available at https://github.com/giorgioroffo/infinite-self-attention.
Acknowledgements
This paper is a preprint; the camera-ready version is expected in September 2026. The author thanks the anonymous reviewers for their constructive feedback during the peer-review process.

### v2

@@ 9 Additional Attention Quality Curves
Together with the summary in the main paper, these four curves confirm that InfSA produces spatially selective, semantically aligned attention across all evaluation protocols.
10 Full Classification Results
Table 3 reports per-method Top-1 accuracy on ImageNet-1K and ImageNet-V2, grouped by attention family. These numbers underlie the accuracy-vs-parameters scatter plots in the main paper (Fig. 6).
@@ 10 Full Classification Results
ViT-B/16 [dosovitskiy2021vit] ’21 86 17.6 81.5 74.1
DeiT-B [touvron2021training] ’21 86 17.5 81.8 71.5
HorNet-B [rao2022hornet] ’22 88 15.5 84.3
SG-Former-B [ren2023sgformer] ’23 78 15.6 84.7
SMT-L [lin2023smt] ’23 81 17.7 84.6
InternImage-B [wang2023internimage] ’23 97 16.0 84.9
GC-ViT-B [hatamizadeh2023gcvit] ’23 90 14.8 85.0
FAT-B5 [fan2023fat] ’23 88 15.1 85.2
STViT-L [chen2023stvit] ’23 95 15.6 85.3
RMT-L [fan2024rmt] ’24 95 18.2 85.5 76.3
@@
SOFT-H [lu2021soft] ’21 87 16.3 83.3
FLatten-S-B [han2023flatten] ’23 89 15.4 83.8
FLatten-C-B [han2023flatten] ’23 75 15.0 84.5
ViG-S [liao2024vig] ’24 23 3.5 83.4
Agent Attn-B [han2024agent] ’24 88 15.4 84.1
MLLA-B [han2024mlla] ’24 96 16.2 85.3 76.7
RAVLT-L [fan2024ravlt] ’24 95 16.0 85.8 76.8
@@
HyenaPixel [spravil2024hyenapixel] ’24 111 25.3 84.9
MambaVision-L [hatamizadeh2025mambavision] ’25 228 34.9 85.0
MambaVision-L2 [hatamizadeh2025mambavision] ’25 242 37.5 85.3
@@
[…]
@@ 11 Full Ablation Results
Key observations. (i) Top-1 accuracy and MoRF-AOC peak at $\gamma{=}0.7$. (ii) LeRF-AUC increases monotonically with $\gamma$ (62.1% $\to$ 65.3%), indicating that deeper path propagation preserves robustness to irrelevant regions. (iii) Latency and convergence remain stable across all settings (35–37 ms train, 11–12 ms infer, 40–45 epochs), confirming negligible overhead from hyper-parameter variation. (iv) Among activations, $|x|$ yields the highest LeRF-AUC (65.5%) at the cost of lower classification accuracy; ReLU offers the best overall trade-off and is adopted as default ($\gamma{=}0.7$, ReLU).
12 Additional Efficiency Analysis
The main paper (Fig. 2(b)) presents the complexity-tier scatter. Here we complement it with a depth-comparison view.
4L vs. 24L depth comparison (Fig. 9). Increasing depth from 4 to 24 layers reduces throughput for all models (e.g., InfViT Linear: 231 $\to$ 40 img/s), yet InfViT Linear retains its ranking advantage in both throughput and energy in both regimes. The $\mathcal{O}(Nd)$ baselines (Linformer, SOFT, Agent Attn) cluster at $10{-}11{\times}$ speed-up, while the $\mathcal{O}(Nd^{2})$ methods (Performer, FLatten, MLLA) reach only $5{-}6{\times}$.
13 Qualitative Analysis of Attention Maps
[…]
@@ 13 Qualitative Analysis of Attention Maps
Semantic localization. InfSA variants (both Pure and Linear) exhibit sharper focus on semantically meaningful object regions (faces, limbs, object centers), whereas softmax attention often highlights diffuse or peripheral areas.
Consistency. Both InfSA variants produce attention distributions that are tighter and more consistent across samples and categories, suggesting higher robustness to irrelevant context. We attribute this to the spectral nature of InfSA: by modeling attention as a power series over token interactions or approximating dominant eigenvectors of the token graph, InfSA captures multi-hop dependencies and structural centrality, leading to localized and informative spatial activations.
14 Reproducibility and Implementation Details
14.1 Latency vs. Resolution
@@ 14 Reproducibility and Implementation Details
14.2 Hardware and Software Environment
@@ 14.4 Data Augmentation
14.5 Model Architecture and FLOPs
@@
14.6 Extreme-Resolution Inference
@@ 15 Discussion and Future Directions
The spectral and diffusion viewpoints rely on standard assumptions (homogeneous operators in the Neumann-series discussion, nonnegativity in Perron–Frobenius analysis). These serve as modelling tools for design motivation; our eigenvector-alignment study (Sect. 8) confirms they hold in practice for trained models.